# Is Your Paper Being Reviewed by an LLM? A New Benchmark Dataset and Approach for Detecting AI Text in Peer Review

### v1

@@ D.2 Comparison of numeric scores assigned by human and AI reviewers
Figure 3: Difference between AI and human scores. For each matched review (aligned by paper ID and recommendation), score differences were computed and displayed as histograms. Scores range from 1 to 4 for all metrics except Confidence, which ranges from 1 to 5. Statistical significance was assessed using a two-sided Wilcoxon signed-rank test, with p-values shown in the legend. This figure includes only NeurIPS2022–2024 and ICLR2024, because they are the onyl conferences that required reviewers to submit these scores in their review templates.
@@ Appendix F Artifact Use Consistent With Intended Use
Appendix G Use of AI Tool
The authors of this paper used Github Co-pilot for coding assistance for analysis.

### v2

@@ D.4 Comparison of numeric scores assigned by human and AI reviewers
Figure S3: Difference between AI and human scores. For each matched review (aligned by paper ID and recommendation), score differences were computed and displayed as histograms. Scores range from 1 to 4 for all metrics except Confidence, which ranges from 1 to 5. Statistical significance was assessed using a two-sided Wilcoxon signed-rank test, with p-values shown in the legend. This figure includes only NeurIPS2022–2024 and ICLR2024, because they are the onyl conferences that required reviewers to submit these scores in their review templates.
D.5 Validation of AI-edited reviews
To simulate varying degrees of LLM-assisted editing, we took human-written reviews and generated AI-edited versions at four levels of modification: Minimum, Moderate, Extensive, and Maximum (see Appendix E.3 for the editing prompts used). To validate that these prompts produce meaningfully distinct levels of change, we computed cosine similarity scores between each AI-edited review and its original human version using text embeddings. As shown in Table S12, similarity scores progressively decrease with greater editing intensity, confirming that the levels are semantically distinguishable.
Table S12: Similarity score of edited reviews and original human-written reviews using different prompts.
Minimum Moderate Extensive Maximum
0.9841 0.9261 0.8616 0.6799
@@ E.3 Editing Prompts
E.4 Archetype Prompts
"Balanced":
You provide fair, balanced, thorough, and constructive feedback, objectively highlighting both the strengths and weaknesses of the paper. You maintain a high standard for research in your decision-making process. However, even if your decision is to reject, you offer helpful suggestions for improvement.}
"Conservative":
You generally prefer established methods and are skeptical of unproven (that is, new or unconventional) approaches. While you maintain high standards and rigor, you are critical of papers presenting new ideas without extensive evidence and thorough validation against established baselines. You place significant emphasis on methodological soundness and are cautious about endorsing innovations that haven’t been rigorously tested.}
"Innovative":
You highly value novelty and bold approaches, often prioritizing novel ideas over methodological perfection. While you maintain high standards, you are willing to overlook minor flaws or incomplete validations and may accept the paper, if the paper introduces a significant new concept or direction. Conversely, you tend to be less enthusiastic about papers that, despite thorough methodology and analysis, offer only incremental improvements, and may recommend rejection for such submissions.}
"Nitpicky":
You are a perfectionist who meticulously examines every aspect of the paper, including minor methodological details, technical nuances, and formatting inconsistencies. Even if a paper presents novel ideas or significant contributions, you may still recommend rejection if you identify a substantial number of minor flaws. Your stringent attention to detail can sometimes overshadow the broader significance of the work in your decision-making process.}
@@ Appendix F Artifact Use Consistent With Intended Use
Appendix G Ethics Statement
Appendix H Use of AI Tool
GitHub Copilot and ChatGPT were used to aid in coding for analysis and in editing text for clarity.